# Aditya Tanna, Pratinav Seth, Mohamed Bouadi Utsav Avaiya, Vinay Kumar Sankarapu Lexsi Labs, India & France

### v2

@@ 6.2.3  Domain-Specific Analysis
XGBoost 6.32 0.7834 0.7669 6.62 0.7958 0.7885
RandomForest 6.38 0.7779 0.7752 7.32 0.8052 0.8001
CatBoost 6.36 0.7784 0.7594 5.82 0.8117 0.8015
LightGBM 5.32 0.7949 0.7614 6.17 0.8095 0.7974
TabICL 5.54 0.7819 0.7696 6.60 0.8125 0.7942
OrionBiX 4.10 0.7893 0.7759 5.39 0.8206 0.8125
OrionMSP 4.50 0.8045 0.7916 4.60 0.8158 0.8047
TabPFN 5.04 0.7984 0.7857 7.17 0.8094 0.7919
Mitra 10.77 0.3935 0.2863 13.67 0.5340 0.4250
ContextTab 8.66 0.6681 0.6129 11.25 0.7430 0.6834
TabDPT 6.86 0.7764 0.7641 8.00 0.8080 0.7960
@@
Medical Domain Performance.
On medical datasets, TabPFN with supervised fine-tuning (SFT) achieves the best overall results (Rank = 1.86, ACC = 0.8094, F1 = 0.7958), followed by its meta-learning variant (Rank = 2.28, ACC = 0.8133, F1 = 0.8000). OrionMSP performs strongly in zero-shot (ACC = 0.8045, F1 = 0.7916) and meta-learning settings (Rank = 2.20, ACC = 0.7744, F1 = 0.7645), but exhibits minor degradation under full SFT. TabDPT (PEFT–SFT) also performs competitively (Rank = 1.96, ACC = 0.7680, F1 = 0.7531), achieving near-optimal results with reduced fine-tuning cost. In contrast, transformer-heavy models such as TabICL and OrionBiX lose up to 7–10 % accuracy under SFT, confirming overfitting risk in small, noisy clinical datasets. Traditional baselines (XGBoost, LightGBM) remain competitive (ACC $\approx$ 0.79–0.80) but trail the best TFMs by 2–3 . Overall, the medical results demonstrate that Bayesian and meta-learning TFMs generalize more reliably under limited data and high noise.
Finance Domain Performance.
Financial datasets are typically larger and more imbalanced, amplifying the value of episodic and parameter-efficient adaptation. OrionMSP under meta-learning attains the best overall rank (2.26) and strong accuracy (ACC = 0.8209, F1 = 0.8089), narrowly surpassing TabPFN (meta, Rank = 3.00, ACC = 0.8220, F1 = 0.8056) and its SFT counterpart (Rank = 2.00, ACC = 0.8222, F1 = 0.8058). TabDPT (PEFT–SFT) maintains competitive accuracy (Rank = 2.82, ACC = 0.7782, F1 = 0.7543), while OrionMSP (PEFT–SFT) achieves the strongest overall efficiency–accuracy balance (Rank = 1.93, ACC = 0.7854, F1 = 0.7272). Transformer models such as TabICL degrade sharply under SFT (ACC drop > 12 %), whereas classical baselines remain robust (ACC $\approx$ 0.81) but underperform in F1 due to poor minority calibration. The finance domain thus emphasizes meta-learning as the most stable strategy for large-scale, imbalanced tabular data.
Discussion.
Across both domains, pretrained TFMs consistently outperform traditional baselines when appropriate fine-tuning strategies are used. TabPFN achieves the best reliability and calibration on small, noisy medical datasets, whereas OrionMSP dominates in large, high-imbalance financial contexts. TabDPT maintains strong performance in both settings through efficient PEFT adaptation, confirming that resource-efficient methods can match full SFT accuracy within 2 %. These trends highlight the need for strategy–domain alignment: Bayesian priors excel under data scarcity, while meta-learning transformers scale better with complex, heterogeneous features.
@@ 6.3  Calibration Evaluation
Episodic meta-learning largely preserves calibration quality. TabPFN achieves the best overall calibration across suites (ECE=0.0271–0.0388, BS=0.118–0.150), improving slightly over zero-shot results. OrionMSP remains competitive (ECE=0.0340–0.0365, BS=0.125–0.166), maintaining stable confidence estimates even after adaptation. In contrast, transformer-heavy models such as TabICL exhibit severe degradation on TabZilla (ECE rises from 0.0369 to 0.2033), showing that full episodic updates can destabilize latent priors when domain variability is high. Interestingly, Mitra improves calibration dramatically under meta-learning (ECE drops from $\approx 0.22$ to 0.030.030.050.05), although its Brier scores remain poor (>0.34>0.34), suggesting partial correction without overall reliability gains.
@@
Discussion
Calibration analysis reveals a clear hierarchy of reliability. TabPFN consistently exhibits the lowest ECE and Brier Scores across all adaptation strategies, validating its Bayesian design and robustness to parameter updates. Meta-learning preserves calibration better than full SFT, confirming that episodic adaptation reinforces pretrained uncertainty priors. OrionMSP achieves strong calibration, especially under zero-shot and meta-learning, while transformer-centric models like TabICL and OrionBiX show pronounced drift once fully fine-tuned. Overall, TFMs maintain lower calibration error than ensemble baselines and degrade far less under fine-tuning, making them more trustworthy for risk-sensitive deployment.
@@ 6.4  Fairness Evaluation
Fairness analysis evaluates whether model predictions exhibit systematic bias across demographic subgroups, complementing accuracy and calibration assessments by addressing equity in high-stakes applications such as lending, healthcare, and criminal justice. Unlike performance metrics that can be computed automatically, fairness requires explicit identification of sensitive attributes (e.g., race, gender, age, income). We use datasets with demographic labels—Adult Census Income, German Credit, and COMPAS Recidivism—to measure bias propagation and mitigation across strategies. Following standard fairness literature, we report Statistical Parity Difference (SPD), Equalized Odds Difference (EOD), and Equalized Opportunity Difference (EOpD). All metrics are absolute differences in the 0–1 range (lower = fairer). Table 9 summarizes fairness–accuracy trade-offs across zero-shot, meta-learning, and fine-tuning strategies.
Zero-shot inference.
Meta-learning fine-tuning.
Episodic adaptation largely preserves fairness while sustaining accuracy. OrionBiX achieves the lowest overall bias (SPD = 0.2021, EOD = 0.1624, EOpD = 0.2010) while maintaining high accuracy (ACC = 0.8743). OrionMSP follows closely (EOD = 0.2798, EOpD = 0.2907), offering the best fairness–accuracy equilibrium across models. TabPFN achieves the highest accuracy (ACC = 0.8733) with slightly higher disparity (SPD $\approx$ 0.31). Mitra remains highly fair but again sacrifices accuracy. Overall, meta-learning fine-tuning maintains fairness levels similar to zero-shot inference while modestly improving accuracy consistency.
Supervised fine-tuning (SFT).
@@ 7.1  Fine-Tuning Strategy Recommendations
  • Zero-shot inference achieves excellent calibration (ECE <<0.04 for top models) and competitive predictive performance for OrionMSP and TabPFN. It is ideal when computational resources are limited or calibration reliability is a priority. Zero-shot OrionMSP achieves strong accuracy on large datasets (ACC = 0.8722–0.8821, ranks 3.263.263.843.84) and top performance on OpenML-CC18 and TabZilla without requiring additional training.
  • Meta-learning fine-tuning provides the best balance between adaptation and generalization. It improves rank stability (OrionMSP: ranks 1.731.732.822.82 vs. 2.902.903.843.84 in zero-shot) while preserving calibration (ECE <<0.06). Meta-learning also excels in high-imbalance financial data (OrionMSP ranks $2.26$, TabPFN ranks $3.00$), confirming its cross-domain robustness. However, instability can occur for some models such as TabICL on TabZilla (ECE increases to $0.2033$), requiring careful validation.
@@ 7.4  Model Selection Guidelines
  • Small datasets (<<1K samples): TabPFN achieves the best results in both zero-shot and meta-learning settings (ACC = 0.8325–0.8336, F1 = 0.8131–0.8256, ranks <2.5<2.5)), effectively capturing few-shot patterns without fine-tuning. Zero-shot OrionMSP offers a robust alternative for rapid deployment (ACC = 0.8232, F1 = 0.8194). Fine-tuning transformer-based models such as TabICL or OrionBiX should be avoided on small datasets due to severe overfitting.
@@ 7.5  Domain-Specific Guidance
Finance domain: OrionMSP under meta-learning achieves the best overall results (Rank = 2.16, ACC = 0.8209, F1 = 0.8128), followed closely by TabPFN (Meta/SFT, Ranks = 2.50–2.89, ACC = 0.8162–0.8784). TabDPT remains competitive (PEFT–SFT, Rank = 2.82) and offers an efficient alternative with only a minor loss in accuracy. These outcomes confirm that episodic and parameter-efficient adaptation strategies are particularly effective for large-scale, class-imbalanced financial data, where they improve both calibration and minority-class performance.
@@ 7.6  Practical Value of Tabular Foundation Models
Key strengths include: (I) Zero-shot adaptability: OrionMSP achieves top-tier zero-shot performance (ACC = 0.8722–0.8821, overall rank $\approx 3.5$) without any task-specific training. (II) Calibration: TabPFN maintains low calibration error (ECE <0.05<0.05) across both zero-shot and fine-tuned regimes, outperforming ensemble baselines whose ECE often exceeds $0.08$. (III) Fairness stability: TFMs achieve moderate statistical parity differences (SPD $\approx 0.28$$0.30$) while sustaining high accuracy (>0.85>0.85), indicating balanced performance across sensitive groups.

### v3

@@ 6.2.3  Domain-Specific Analysis
XGBoost 7.32 0.7834 0.7669 7.62 0.7958 0.7885
RandomForest 7.30 0.7779 0.7752 8.46 0.8052 0.8001
CatBoost 7.48 0.7784 0.7594 6.75 0.8117 0.8015
LightGBM 6.36 0.7949 0.7614 7.32 0.8095 0.7974
TabICL 6.66 0.7819 0.7696 7.82 0.8125 0.7942
OrionBiX 4.98 0.7893 0.7759 6.46 0.8206 0.8125
OrionMSP 5.31 0.8045 0.7916 5.25 0.8158 0.8047
TabPFN-2.0 5.82 0.7984 0.7857 8.21 0.8094 0.7919
TabPFN-2.5 5.04 0.7965 0.7835 4.57 0.8169 0.8000
Mitra 12.00 0.3935 0.2863 15.14 0.5340 0.4250
ContextTab 6.40 0.8003 0.7881 10.85 0.8000 0.7830
TabDPT 7.96 0.7764 0.7641 9.46 0.8080 0.7960
@@
Medical Domain Performance : On medical datasets, TabPFN with supervised fine-tuning (SFT) achieves the best overall results (Rank = 1.86, ACC = 0.8094, F1 = 0.7958), followed by its meta-learning variant (Rank = 2.28, ACC = 0.8133, F1 = 0.8000). OrionMSP performs strongly in zero-shot (ACC = 0.8045, F1 = 0.7916) and meta-learning settings (Rank = 2.20, ACC = 0.7744, F1 = 0.7645), but exhibits minor degradation under full SFT. TabDPT (PEFT–SFT) also performs competitively (Rank = 1.96, ACC = 0.7680, F1 = 0.7531), achieving near-optimal results with reduced fine-tuning cost. In contrast, models such as TabICL and OrionBiX lose up to 7–10 % accuracy under SFT, confirming overfitting risk in small, noisy clinical datasets. Traditional baselines (XGBoost, LightGBM) remain competitive (ACC $\approx$ 0.79–0.80) but trail the best TFMs by 2–3 . Overall, the medical results demonstrate that meta-learning TFMs generalize more reliably under limited data and high noise.
Finance Domain Performance : Financial datasets are typically larger and more imbalanced, amplifying the value of episodic and parameter-efficient adaptation. OrionMSP under meta-learning attains the best overall rank (2.26) and strong accuracy (ACC = 0.8209, F1 = 0.8089), narrowly surpassing TabPFN (meta, Rank = 3.00, ACC = 0.8220, F1 = 0.8056) and its SFT counterpart (Rank = 2.00, ACC = 0.8222, F1 = 0.8058). TabDPT (PEFT–SFT) maintains competitive accuracy (Rank = 2.82, ACC = 0.7782, F1 = 0.7543), while OrionMSP (PEFT–SFT) achieves the strongest overall efficiency–accuracy balance (Rank = 1.93, ACC = 0.7854, F1 = 0.7272). Transformer models such as TabICL degrade sharply under SFT (ACC drop > 12 %), whereas classical baselines remain robust (ACC $\approx$ 0.81) but underperform in F1 due to poor minority calibration. The finance domain thus emphasizes meta-learning as the most stable strategy for large-scale, imbalanced tabular data.
Discussion : Across both domains, pretrained TFMs consistently outperform traditional baselines when appropriate fine-tuning strategies are used. TabPFN achieves the best reliability and calibration on small, noisy medical datasets, whereas OrionMSP dominates in large, high-imbalance financial contexts. TabDPT maintains strong performance in both settings through efficient PEFT adaptation, confirming that resource-efficient methods can match full SFT accuracy within 2 %. These trends highlight the need for strategy–domain alignment: Bayesian priors excel under data scarcity, while meta-learning transformers scale better with complex, heterogeneous features.
@@ 6.3  Calibration Evaluation
Episodic meta-learning largely preserves calibration quality. TabPFN achieves the best overall calibration across suites (ECE=0.0271–0.0388, BS=0.118–0.150), improving slightly over zero-shot results. OrionMSP remains competitive (ECE=0.0340–0.0520, BS=0.125–0.166), maintaining stable confidence estimates even after adaptation. In contrast, transformer-heavy models such as TabICL exhibit severe degradation on TabZilla (ECE rises from 0.0369 to 0.2033), showing that full episodic updates can destabilize latent priors when domain variability is high. Interestingly, Mitra improves calibration dramatically under meta-learning (ECE drops from $\approx 0.22$ to 0.030.030.050.05), although its Brier scores remain poor (>0.34>0.34), suggesting partial correction without overall reliability gains.
@@
Discussion : Calibration analysis reveals a clear hierarchy of reliability. TabPFN consistently maintains among the lowest ECE and Brier scores across all adaptation strategies (ECE 0.027–0.047, BS 0.12–0.16), validating its Bayesian design and robustness to parameter updates. Meta-learning preserves calibration better than full SFT, confirming that episodic adaptation reinforces pretrained uncertainty priors. OrionMSP achieves strong calibration, especially under zero-shot and meta-learning, while transformer-centric models like TabICL and OrionBiX show pronounced drift once fully fine-tuned. Overall, TFMs maintain lower calibration error than ensemble baselines and degrade far less under fine-tuning, making them more trustworthy for risk-sensitive deployment.
@@ 6.4  Fairness Evaluation
Fairness analysis evaluates whether model predictions exhibit systematic bias across demographic subgroups, complementing accuracy and calibration assessments by addressing equity in high-stakes applications such as lending, healthcare, and criminal justice. Unlike performance metrics that can be computed automatically, fairness requires explicit identification of sensitive attributes (e.g., race, gender). We use datasets with demographic labels—Adult Census Income, German Credit, and COMPAS Recidivism—to measure bias propagation and mitigation across strategies. Following standard fairness literature, we report Statistical Parity Difference (SPD), Equalized Odds Difference (EOD), and Equalized Opportunity Difference (EOpD). All metrics are absolute differences in the 0–1 range (lower = fairer). Table 9 summarizes fairness–accuracy trade-offs across zero-shot, meta-learning, and fine-tuning strategies.
Zero-shot inference.
Meta-learning fine-tuning.
Episodic adaptation largely preserves fairness while sustaining accuracy. OrionBiX achieves the lowest overall bias (SPD = 0.2021, EOD = 0.1624, EOpD = 0.2010) while maintaining high accuracy (ACC = 0.8743). OrionMSP follows closely (EOD = 0.2798, EOpD = 0.2907),a close second with slightly higher bias but similar accuracy.
TabPFN achieves the highest accuracy (ACC = 0.8733) with slightly higher disparity (SPD $\approx$ 0.31). Mitra remains highly fair but again sacrifices accuracy. Overall, meta-learning fine-tuning maintains fairness levels similar to zero-shot inference while modestly improving accuracy consistency.
Supervised fine-tuning (SFT).
@@ 7.1  Fine-Tuning Strategy Recommendations
  • Zero-shot inference achieves excellent calibration (ECE <<0.04 for top models) and competitive predictive performance. It is ideal when computational resources are limited or calibration reliability is a priority. TabPFN-2.5 attains the highest overall zero-shot standing (Rank = 4.13), establishing a new state-of-the-art for inference-only tabular tasks. OrionMSP follows closely, achieving strong accuracy on large datasets (ACC = 0.8722–0.8821) and ranking among the top configurations on OpenML-CC18 and TabZilla without requiring additional training.
  • Meta-learning fine-tuning provides the best balance between adaptation and generalization. It improves rank stability for models such as OrionMSP (aggregate rank $2.26$ in Table 3) while preserving calibration (ECE <<0.06). Meta-learning also excels in high-imbalance settings (OrionMSP rank $2.16$, TabPFN rank $2.50$ in Table 6), confirming its cross-domain robustness. However, instability can occur for some models such as TabICL on TabZilla (ECE increases to $0.2033$), requiring careful validation.
@@ 7.4  Model Selection Guidelines
  • On small datasets (<<1K samples), TabPFN-2.5 achieves the highest zero-shot performance (ACC = 0.8349, F1 = 0.8194), slightly ahead of TabDPT (0.8333/0.8271) and boosted-tree baselines. Under meta-learning, TabPFN-2.0 also leads (0.8336/0.8256), while OrionMSP (PEFT–Meta) provides strong performance at lower fine-tuning cost (0.8275/0.8213). Fine-tuning TabICL or OrionBiX on small datasets still causes sharp degradation (ACC drops $15\%$), revealing overfitting risk.
@@ 7.5  Domain-Specific Guidance
Finance domain: OrionMSP under meta-learning achieves strong results (Rank = 2.26, ACC = 0.8209, F1 = 0.8089), while TabPFN performs exceptionally well in supervised and meta-learning modes (Ranks = 2.00--3.00, ACC = 0.8220--0.8222). TabDPT remains competitive (PEFT--SFT, Rank = 2.82) and offers an efficient alternative with only a minor loss in accuracy. These outcomes confirm that episodic and parameter-efficient adaptation strategies are particularly effective for large-scale, class-imbalanced financial data, where they improve both calibration and minority-class performance.
@@ 7.6  Practical Value of Tabular Foundation Models
Key strengths include: (I) ZERO-SHOT ADAPTABILITY: The TabPFN family and OrionMSP provide top-tier zero-shot performance. TabPFN-2.5 attains the best aggregate zero-shot rank (4.13) with ACC $\approx\textsc{0.85}$0.88, while OrionMSP remains highly competitive (rank 4.61, ACC = 0.84550.8821). Both outperform boosted-tree baselines by roughly 1.5–2.5 percentage points in accuracy without any task-specific training. (II) Calibration: TabPFN maintains low calibration error (ECE <0.05<0.05) across both zero-shot and fine-tuned regimes, outperforming ensemble baselines whose ECE often exceeds $0.08$. (III) Fairness stability: TFMs achieve moderate statistical parity differences (SPD $\approx 0.28$$0.30$) while sustaining high accuracy (>0.85>0.85), indicating balanced performance across sensitive groups.